# Gaussian Adaptive Attention is All You Need: Robust Contextual Representations Across Multiple Modalities

### 20 Jan 2024 v1

@@ Abstract
We propose the Multi-Head Gaussian Adaptive Attention Mechanism (GAAM), a novel probabilistic attention framework, and the Gaussian Adaptive Transformer (GAT), designed to enhance information aggregation across multiple modalities, including Speech, Text and Vision. GAAM integrates learnable mean and variance into its attention mechanism, implemented in a Multi-Headed framework enabling it to collectively model any Probability Distribution for dynamic recalibration of feature significance. This method demonstrates significant improvements, especially with highly non-stationary data, surpassing the state-of-the-art attention techniques in model performance (up to approximately +20% in accuracy) by identifying key elements within the feature space. GAAM’s compatibility with dot-product-based attention models and relatively low number of parameters showcases its adaptability and potential to boost existing attention frameworks. Empirically, GAAM exhibits superior adaptability and efficacy across a diverse range of tasks, including emotion recognition in speech, image classification, and text classification, thereby establishing its robustness and versatility in handling multi-modal data. Furthermore, we introduce the Importance Factor (IF), a new learning-based metric that enhances the explainability of models trained with GAAM-based methods. Overall, GAAM represents an advancement towards development of better performing and more explainable attention models across multiple modalities.

### 25 Jan 2024 v2

@@ Abstract
We propose the Multi-Head Gaussian Adaptive Attention Mechanism (GAAM), a novel probabilistic attention framework, and the Gaussian Adaptive Transformer (GAT), designed to enhance information aggregation across multiple modalities, including Speech, Text and Vision. GAAM integrates learnable mean and variance into its attention mechanism, implemented in a Multi-Headed framework enabling it to collectively model any Probability Distribution for dynamic recalibration of feature significance. This method demonstrates significant improvements, especially with highly non-stationary data, surpassing the state-of-the-art attention techniques in model performance (up to approximately +20% in accuracy) by identifying key elements within the feature space. GAAM’s compatibility with dot-product-based attention models and relatively low number of parameters showcases its adaptability and potential to boost existing attention frameworks. Empirically, GAAM exhibits superior adaptability and efficacy across a diverse range of tasks, including emotion recognition in speech, image classification, and text classification, thereby establishing its robustness and versatility in handling multi-modal data. Furthermore, we introduce the Importance Factor (IF), a new learning-based metric that enhances the explainability of models trained with GAAM-based methods. Overall, GAAM represents an advancement towards development of better performing and more explainable attention models across multiple modalitiesCode: https://github.com/gioannides/Gaussian-Adaptive-Attention.
@@ 5 Conclusion
Appendix
Extended Results
In order to address the lack of fold-based performance evaluation, similar to what is available when using the IEMOCAP dataset, we conduct additional runs on the CIFAR100 and AG News datasets. This approach is necessary to ensure a robust validation of our methods, as the original datasets do not provide the opportunity for a fold-based analysis. We present the results of these additional runs, offering a more comprehensive view of the performance consistency and effectiveness of GAAM-based methods. In Tables 8 and 9, the first run (i.e. R1) depicts the original results from Tables 4 and 4, while the rest of the runs depict the extended results.
Method R1 R2 R3 R4 R5 Avg.
MHA 0.604 0.619 0.621 0.620 0.621 0.617
MHA $\rightarrow$ BN 0.630 0.671 0.695 0.639 0.670 0.661
GQGAAM 0.800 0.801 0.801 0.806 0.800 0.801
GAAMv1 0.799 0.802 0.802 0.807 0.807 0.803
SGA 0.802 0.807 0.807 0.801 0.808 0.805
bi-SGA 0.802 0.807 0.807 0.801 0.808 0.805
GAAMv2 0.802 0.804 0.810 0.803 0.810 0.806
Table 8: CIFAR100 5 Run validation (R1-R5) using BEiT-Large as Encoder.
For CIFAR100, as demonstrated in Table 8, GAAMv2 overall outperforms other methods, with an average accuracy rate of 0.806. This performance is followed by SGA and bi-SGA, both yielding an average accuracy rate of 0.805. Notably, all GAAM-based methods (GQGAAM, GAAMv1, GAAMv2) significantly surpass the MHA and MHA $\rightarrow$ BN models, underlining the superiority of GAAM frameworks in handling the Vision modality.
Method R1 R2 R3 Avg.
GAAMv2 0.944 0.945 0.946 0.945
MHA $\rightarrow$ BN 0.945 0.945 0.947 0.946
MHA 0.944 0.945 0.948 0.946
SGA 0.945 0.945 0.947 0.946
GAAMv1 0.945 0.945 0.947 0.946
bi-SGA 0.947 0.945 0.947 0.946
GQGAAM 0.948 0.949 0.949 0.949
Table 9: AG News 5 Run validation (R1-R5) using Llama2-13B as Encoder.
Similarly, in the AG News dataset, as shown in Table 9, GQGAAM emerges as the top performer with an average accuracy rate of 0.949, outperforming other methodologies. This consistency in high performance across different runs for GAAM-based methods, compared to MHA and MHA $\rightarrow$ BN, as well as (bi-)SGA methods, validates the hypothesis of GAAM’s effectiveness in the Text modality context.
Gaussian Adaptive Attention Mechanism (GAAM)
In this section, we delve into a more comprehensive analysis of the Gaussian Adaptive Attention Mechanism (GAAM), elucidating its operational intricacies and the pivotal role it plays in enhancing model performance. The GAAM operates on the principle of dynamically adjusting focus on different segments of the input tensor, utilizing learned statistical properties. This adaptive mechanism allows for a nuanced and context-sensitive interpretation of complex data modalities, leading to more accurate and efficient model outcomes. The algorithm for GAAM can be found in listing 1. The algorithm of the Multi-Headed framework and how it utilizes GAAM can be found in listing 2.
Algorithm 1 Gaussian Adaptive Attention Mechanism (GAAM)
0:  $x$ (input tensor), normDimSize, normAxis, c, eps
0:  Attention-modified tensor
1:  Initialize $\mathbf{c}$ to a tensor of size $(1,\text{normDimSize})$ filled with $c$
2:  Initialize $\mathbf{meanOffset}$ to a tensor of size $(1,\text{normDimSize})$ filled with zeros
3:  for each batch in $x$ do
4:     $\mathbf{mean}\leftarrow\text{mean}(x,\text{dim}=\text{normAxis})$
5:     $\mathbf{var}\leftarrow\text{mean}(x^{2},\text{dim}=\text{normAxis})-\mathbf{mean}^{2}$
6:     $\mathbf{var}\leftarrow|\mathbf{var}|+10^{-8}$
7:     $\mathbf{adjustedMean}\leftarrow\mathbf{mean}+\mathbf{meanOffset}$
8:     $\mathbf{yNorm}\leftarrow(x-\mathbf{adjustedMean})/\sqrt{\mathbf{var}+10^{-5}}$
9:     $\mathbf{yTransform}\leftarrow\exp(-(\mathbf{yNorm}^{2}/(2\cdot\mathbf{c})))$
10:     $x\leftarrow x\cdot\mathbf{yTransform}$
11:  end for
12:  return  $x$
Parameters and their Roles
  • $x$ (Input Tensor): The data on which the attention mechanism is applied.
  • normDimSize: Specifies the size of the dimension for normalization.
  • normAxis: Determines the axis of $x$ along which the mean and variance are computed.
  • $c$: A learnable parameter controlling the scale of Gaussian normalization.
  • clear numbers: A small constant for numerical stability.
Learning Aspects:
  • $\mathbf{c}$ and $\mathbf{meanOffset}$ are learnable parameters adjusting the normalization’s impact and bias.
Operational Steps:
  1. 1.
    Compute mean and variance along normAxis.
  2. 2.
    Apply Gaussian normalization to each feature, re-weighted by learned parameters.
  3. 3.
    The output is the input tensor modulated by the learned attention.
Multi-Head Gaussian Adaptive Attention
Algorithm 2 Multi-Head Gaussian Adaptive Attention
0:  $x$ (input tensor), normDimSize, numHeads, normAxis, c
0:  Concatenated attention output tensor
1:  Initialize an array $\mathbf{attentionHeads}$ of size numHeads
2:  for $i=1$ to numHeads do
3:     $\mathbf{attentionHeads}[i]\leftarrow\text{GAAM}($normDimSize, normAxis, c, eps)normal-)))
4:  end for
5:  Initialize an empty list $\mathbf{outputs}$
6:  for each $\mathbf{head}$ in $\mathbf{attentionHeads}$ do
7:     $\mathbf{headOutput}\leftarrow\text{apply}(\mathbf{head},x)$
8:     Append $\mathbf{headOutput}$ to $\mathbf{outputs}$
9:  end for
10:  $\mathbf{output}\leftarrow\text{concatenate}(\mathbf{outputs},\text{dim}=\text{normAxis})$
11:  return  $\mathbf{output}$
GAAM applies multiple instances of GAAM in parallel, allowing simultaneous focus on different parts or aspects of the data.
Parameters and their Roles in Multi-Head GAAM:
  • $x$ (Input Tensor): The data for multi-head attention application.
  • normDimSize: Size of the normalization dimension.
  • numHeads: Number of parallel GAAM mechanisms.
  • normAxis: The axis for normalization.
  • $c$: The scale parameter for Gaussian normalization, shared across heads.
Learning Aspects:
  • Each head in $\mathbf{attentionHeads}$ applies GAAM independently.
Operational Steps:
  1. 1.
    Initialize and apply multiple GAAM instances to the input.
  2. 2.
    Concatenate the outputs from each head along normAxis.
  3. 3.
    The result is a combined attention output, integrating multiple focus areas.